OpenStudy (anonymous):

Jackson goes to the gym 0, 2, or 3 days in a week, depending on work demands at the office. The probability that he goes 0 days is 0.1, the probability that he goes 2 days is 0.65, and the probability that he goes 3 days is 0.25. What is the expected value of the number of days per week for which Jackson goes to the gym?

OpenStudy (anonymous):

\[0\times .1+2\times .65+3\times .25\] is the expected value you can ignore the first term